A 127127‑turn circular coil of radius 3.55 cm3.55 cm and negligible resistance is immersed in a uniform magnetic field that is perpendicular to the plane of the coil. The coil is connected to a 10.9 Ω10.9 Ω resistor to create a closed circuit. During a time interval of 0.161 s,0.161 s, the magnetic field strength decreases uniformly from 0.657 T0.657 T to zero. Find the energy ?E in millijoules that is dissipated in the resistor during this time interval.
Answer #1

Area, A = pi x 0.0355^2 = 0.004 m^2

Emf, E = NA dB/dt

E = 127 x 0.004 x 0.657/0.161 = 2.05 V

Current, I = E/R = 2.05/10.9 = 0.188 A

Energy, e = I^2 Rt = 0.188^2 x 10.9 x 0.161 = 62.2 mJ

Comment in case any doubt please rate my answer....